rssLink RSS for all categories
 
icon_red
icon_green
icon_green
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_orange
icon_green
icon_orange
icon_green
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_red
 

FS#31367 — LNS BOR

Attached to Project— xDSL / FTTH
Incident
Backend / Core
CLOSED
100%
We lost BGP session with FT on our LNS Bordeaux, access should be switched to the other LNS in Paris.
Date:  Friday, 27 April 2018, 15:23PM
Reason for closing:  Done
Additional comments about closing:  Waiting for intervention, no customer impact as far as we can see.
Comment by OVH - Thursday, 26 April 2018, 15:42PM

The issue come from problem the black fiber. We are investigating.


Comment by OVH - Thursday, 26 April 2018, 16:56PM

We also have an electrical issue. We plan a maintenance that will break FT collect access during it.

Expect about 5minute downtime.


Comment by OVH - Friday, 27 April 2018, 00:05AM

starting the operation.


Comment by OVH - Friday, 27 April 2018, 01:23AM

power problem fixed.


Comment by OVH - Friday, 27 April 2018, 09:37AM

Fiber still down, we are waiting for an intervention from the fiber provider.


Comment by OVH - Friday, 27 April 2018, 15:22PM

No impact on customer.
Intervention is planned on monday.